OpenStudy (anonymous):

The length of a rectangle is 6 feet longer than its width. If the area is 50 square feet, find the length and width of the rectangle.

OpenStudy (mathstudent55):

Let L = length, and W = width L = W + 6 Area = LW = 50 (W + 6)W = 50 Solve for W, then find L using L = W + 6

OpenStudy (anonymous):

how do you solve for W?

OpenStudy (mathstudent55):

(W + 6)W = 50 Distribute the W W^2 + 6W = 50 W^2 + 6W - 50 = 0 Now solve the quadratic equation by either factoring, or by the quadratic formula.

OpenStudy (anonymous):

which one will work the best?

OpenStudy (mathstudent55):

This is not factorable, so skip factoring and go straight to the quadratic formula.